# BigTokDetect: A Clinically-Informed Vision–Language Modeling Framework for Detecting Pro-Bigorexia Videos on TikTok

### v2

@@ 5.3 Input Features
To capture the full range of pro-bigorexia signals, we extract four complementary input features: Visual: For models that natively accept video (e.g., Gemini-2.5-Flash, Qwen2.5VL, InternVL3), we provide raw video. For frame-based models (e.g., Claude-Sonnet-4, GPT-4.1, open-source VLMs), we sample four equally spaced frames per video due to API costs and token-length constraints; ablation studies (Section 5.7) show additional frames yield minimal gains. Audio: We apply Google’s YAMNet Hershey et al. (2017) to filter non-speech sounds (music, ambient noise), then transcribe "Speech" segments using OpenAI Whisper Radford et al. (2022). On-Screen Text: We use Gemini-2.5-Flash to extract overlaid text—captions and annotations—since many TikTok creators rely on on-screen text for key messages. Caption: We include the original TikTok description (user caption and hashtags) from the video metadata.
@@
Claude-Sonnet-4 Zero-shot 0.829 0.832 0.829 0.827 0.670 0.679 0.636 0.640
@@
Gemini-2.5-Flash Zero-shot 0.776 0.782 0.776 0.775 0.666 0.665 0.616 0.621
@@
Qwen2.5-VL-7B Zero-shot 0.539 0.744 0.539 0.526 0.383 0.554 0.327 0.306
@@
Qwen2.5-VL-32B Zero-shot 0.733 0.788 0.733 0.742 0.556 0.637 0.495 0.513
@@
InternVL3-8B Zero-shot 0.614 0.734 0.614 0.635 0.517 0.582 0.464 0.476
@@
InternVL3-38B Zero-shot 0.784 0.797 0.784 0.785 0.649 0.665 0.603 0.608
@@
Table 2: Classification Macro Metrics for Primary Category and Subcategory by Model and Training Method. Highest values in each column are in green, second-highest in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
5.3.1 Training and Inference Paradigms
@@
We create prompts for zero-shot classification by incorporating taxonomy definitions of the primary and subcategories and instructing the model to select the appropriate category as the label. The full zeroshot prompt templates for Tasks 1 and 2 are shown in Figure 6 and 7, Appendix D.2. We evaluate the models on balanced test sets of 588 examples for primary classification (Task 1) and 466 examples for subcategory classification (Task 2), with each class evenly represented (see Appendix C)
@@
For few-shot experiments, we sample a fixed set of in-context examples from the training data: two videos per primary category (12 examples total) for Task 1 and one video per subcategory (16 examples) for Task 2 (prompt details are in Figures 8 and 9, Appendix). Each example is presented with its full multimodal features: video frames, audio transcript, on-screen text, caption, and the corresponding label. We evaluate the VLMs on a balanced test set of 462 examples (Appendix C).
@@
Due to cost constraints, we limit finetuning to open-source VLMs. We instruction-tune various model variants (up to 72B) on our annotated datasets. The batch size is 8, and the learning rate is 5e-5. The inference hyperparameters are similar to zero-shot and few-shot prompting (temperature = 0.1 and top_p = 0.9).
@@ 5.4 Pro-Bigorexia Classification Results
Primary Category Detection
Commercial API-based models continue to lead in broad-category detection: Claude-Sonnet-4 (zero-shot) attains the highest accuracy (0.829), macro precision (0.832), macro recall (0.829), and macro F1 (0.827), with its few-shot variant ranking a close second. Although fine-tuned open-source models (e.g., Qwen2.5-VL-32B and InternVL3 variants) gain 10–20 points over their zero-shot baselines, none yet surpass the commercial zero-shot benchmark. These results confirm that large-scale pretraining still confers an advantage for general category classification, even as domain-specific tuning significantly narrows the gap.
Subcategory Detection
In contrast, fine-tuning proves decisive for fine-grained subcategory tasks. InternVL3-8B (finetuned) achieves the top subcategory accuracy (0.690), recall (0.679), and F1 (0.675), while Qwen2.5-VL-7B (finetuned) leads in precision (0.687) and ranks second in other metrics. Commercial zero- and few-shot methods plateau around 0.670 accuracy and underperform relative to these tuned open-source variants. This demonstrates that parameter adaptation is essential to capture the nuanced distinctions of harmful content subtypes.
Prompting vs. Parameter-Tuning
@@
Few-shot prompting yields modest boosts in primary-category performance—for instance, GPT-4.1 improves from 0.796 to 0.813 accuracy—highlighting the continued value of emergent prompting abilities. However, in subcategory detection, few-shot gains are minimal, and zero-shot performance remains low. By contrast, parameter-efficient fine-tuning delivers substantial improvements across both tasks, underscoring that direct model adaptation is the most effective strategy for specialized harmful content detection, particularly for granular subcategory classification.
Model Scale and Architecture
Larger Qwen2.5-VL-32B boosts zero/few-shot primary category performance, but mid-scale InternVL3-8B (built on the Qwen backbone) achieves the strongest fine-tuning gains—surpassing its 38B variant on subcategory tasks. This underscores that, beyond raw parameter count, architecture critically shapes fine-grained detection performance.
@@ 5.5 Severity Score Prediction
Claude-Sonnet-4 0.679 0.675
GPT-4.1 0.690 0.690
Gemini-2.5-Flash 0.690 0.690
InternVL3-8B 0.805 0.474
InternVL3-38B 0.701 0.607
Qwen2.5-VL-7B 0.794 0.484
Qwen2.5-VL-32B 0.688 0.601
Qwen2.5-VL-72B 0.653 0.638
Table 3: Zero-shot Severity Prediction Metrics by Model (highest in green, 2nd highest in blue).
To assess the ability of VLMs to predict harm severity, we conducted zero-shot prompting experiments across multiple models. Results reveal that models typically misestimate severity by approximately half a scale point, indicating reasonable but imperfect alignment with expert clinical judgment. Notably, severity prediction proves more challenging than categorical classification, with all models showing only moderate correlation with ground truth ratings. While these findings demonstrate baseline competence in severity estimation, substantial improvement is needed for reliable moderation deployment.
@@ 5.6 Error Analysis
Cross-model error analysis exposes fundamental challenges in characterizing pro-bigorexia content. For instance, models consistently misclassify exercise-related content as “Irrelevant” (6.1–14.3%) or confuse “Relationship to Body” with “Relationship to Exercise” (10.2–17.3%), mirroring annotator disagreements (Figures 3 and 4, Appendix B.5). Videos demonstrating workout sessions simultaneously touch on exercise, body display, dieting, lifestyle, and motivations that single-label classification cannot capture.
Classification consistently confuses “Hormone Therapy” and “Anabolic Steroids” (11.8–17.6% bidirectional misclassification), a clinically critical distinction. Models also frequently misclassify “Supplement Abuse” as “Relationship to Body” (8.2–22.4%), suggesting difficulty in recognizing harmful supplement messaging. This likely stems from creators using coded language (e.g., “tren”, “stack”), requiring domain expertise that general-purpose models lack. Additionally, models show systematic bias toward predicting “Muscularity Self-objectification” in subtype classification, with excessive exercise and toxic motivation content frequently misclassified into this category (14.7–47.1%). These systematic errors highlight a fundamental challenge: effective pro-bigorexia detection demands not just multimodal capabilities, but clinical and social knowledge to navigate the blurred boundaries between fitness content and promoting unhealthy behaviors.
@@ 5.7.1 Input Features: Text vs Video
Gemini-2.5-Flash Audio 0.700 0.423 0.424
@@
Table 4: Ablation results for task 1 using a single modality input features (text or video). Audio modality refers to the transcript of the video; Caption refers to the video description; OCR is text within images, and Text modality refers to Caption + OCR. Highest values in each column are in green, second-best in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
@@ 5.7.2 Number of Frames per Video
Table 5: Frame-count ablation for GPT-4.1 and InternVL3-38B. Pm, Rm, and F1m are macro precision, macro recall, and macro F1.
@@ 6 Conclusion
We present BigTok and BigTokDetect, a clinically-informed dataset and vision–language framework for identifying pro-bigorexia content on TikTok. By curating over 2,200 expert-annotated videos across five primary categories and fine-grained subcategories, we enable systematic study of muscle dysmorphic signals that evade traditional text-based moderation. Our fine-tuned VLMs achieve 82.9% accuracy on primary category classification and 69.0% on subcategory detection, and ablation studies confirm that multimodal fusion of visual, audio, and textual inputs yields a 5–10% boost over unimodal baselines.
Despite these advances, challenges remain in scaling to continually evolving platforms and detecting emerging coded language. Future work expands BigTok to include cross-platform content (e.g., Instagram, YouTube Shorts), incorporates temporal dynamics of video sequences, and integrates user-level metadata for context-aware moderation. We also collaborate with clinical practitioners to validate model outputs in real-world settings and explore fairness considerations across demographic groups. By releasing our dataset, code, and model checkpoints, we aim to catalyze further research on robust, ethically grounded multimodal moderation in specialized mental health domains.
@@
This research addresses harmful content classification, which raises important ethical considerations. To protect annotators’ well-being when reviewing potentially disturbing pro-bigorexia content, we exclusively recruited clinical experts—licensed psychologists, psychiatrists, and doctoral candidates specializing in eating disorders—who possess the professional training necessary to safely engage with this sensitive material. Their clinical expertise provides essential psychological safeguards while ensuring high-quality annotations. While our work aims to support mental health research, we acknowledge that automated detection systems risk censorship or discrimination against legitimate fitness communities. We took precautions, including content anonymization and clinical expert consultation, focusing on detection research rather than deployment recommendations. Our work is intended for research purposes and should not be used for clinical diagnosis of bigorexia without proper clinical oversight. All procedures were conducted in accordance with institutional review board guidelines.
We acknowledge the use of AI language models to assist with writing and editing portions of this manuscript. All AI-generated content was reviewed, verified, and edited by the authors. The research design, data collection, analysis, and primary contributions remain entirely the work of the human authors.

### v3

@@ 5.3 Input Features
To capture the full range of pro‑bigorexia signals, we extract four complementary input features: Visual: We adapt visual inputs to model architectures. For Gemini-2.5-Flash and Qwen2.5-VL, which support native video processing, we provide the raw video stream. For InternVL3, GPT-4.1, and Claude-Sonnet-4, we treat video as a sequential image task. We extract 4 equally spaced frames per video (unless otherwise noted in ablation studies) to balance temporal coverage with context-window constraints. Audio: We classify audio using YAMNet Hershey et al. (2017) and transcribe detected speech via Whisper Radford et al. (2022). To prevent bias, we retain all videos regardless of audio type: 40.2% contain intelligible speech, while the remaining 59.8% (music or ambient noise) are assigned empty transcripts. On‑Screen Text: We use Gemini‑2.5‑Flash to extract overlaid text, creators’ captions and annotations, since many TikTok creators rely on on‑screen text for key messages. Caption: We include the original TikTok description (user caption and hashtags) from the video metadata.
@@
Claude‐Sonnet‐4 Zero-shot 0.829 0.832 0.829 0.827 0.670 0.679 0.636 0.640
@@
Gemini‐2.5‐Flash Zero-shot 0.805 0.807 0.805 0.805 0.666 0.665 0.616 0.621
@@
Qwen2.5‐VL‐7B Zero-shot 0.539 0.744 0.539 0.526 0.383 0.554 0.327 0.306
@@
Qwen2.5‐VL‐32B Zero-shot 0.733 0.788 0.733 0.742 0.556 0.637 0.495 0.513
@@
InternVL3‐8B Zero-shot 0.614 0.734 0.614 0.635 0.517 0.582 0.464 0.476
@@
InternVL3‐38B Zero-shot 0.784 0.797 0.784 0.785 0.649 0.665 0.603 0.608
@@
Table 2: Classification Macro Metrics for Primary Category and Subcategory by Model and Training Method. Highest values in each column are in green, second‑highest in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
5.4 Training and Inference Paradigms
@@
We create prompts for zero‐shot classification by incorporating taxonomy definitions of the primary and subcategories and instructing the model to select the appropriate category as the label. The full zero-shot prompt templates for Tasks 1 and 2 are shown in Figure 8 and 9, Appendix D.2. We evaluate the models on balanced test sets of 588 examples for primary classification (Task 1) and 466 examples for subcategory classification (Task 2), with each class evenly represented (Appendix C)
@@
For few‐shot experiments, we sample a fixed set of in‐context examples from the training data: two videos per primary category (12 examples total) for Task 1 and one video per subcategory (16 examples) for Task 2 (prompt details are in Figures 10 and 11, Appendix). Each example is presented with its full multimodal features: video frames, audio transcript, on‐screen text, caption, and the corresponding label. We evaluate the VLMs on a balanced test set of 462 examples (Appendix C).
@@
Due to cost constraints, we limit finetuning to open‐source VLMs. We instruction‐tune various model variants (up to 72B) on our annotated datasets. The batch size is 8, and the learning rate is 5e-5. The inference hyperparameters are similar to zero-shot and few-shot prompting (temperature = 0.1 and top_p = 0.9).
@@ 5.5 Pro-Bigorexia Classification Results
Table 2 shows that while large commercial models like Claude-Sonnet-4 achieve the highest zero-shot performance on broad primary categories, parameter-efficient finetuning allows smaller open-source models (e.g., InternVL3-8B) to surpass them on fine-grained subcategory detection by capturing domain-specific clinical nuances.
Primary vs. Subcategory Detection
Commercial API-based models continue to lead in broad-category detection: Claude-Sonnet-4 (zero-shot) attains the highest accuracy of 0.829 and F1 of 0.827. Although few-shot open-source models (e.g., InternVL3-38B) reach competitive levels (0.806 accuracy), none yet surpass the commercial zero-shot benchmark on primary categories. However, the trend reverses for fine-grained subcategory detection tasks. InternVL3-8B (finetuned) achieves the top subcategory F1 of 0.675, outperforming commercial methods, which plateau around 0.640 F1. This demonstrates that while large-scale pretraining confers an advantage for general classification, parameter adaptation better captures nuanced subtype distinctions.
Impact of Model Scale
Our experiments reveal a divergence in finetuning effectiveness relative to model scale. While finetuning dramatically boosts smaller models (7B–8B), for instance, raising Qwen2.5-VL-7B subcategory accuracy from 0.383 (zero-shot) to 0.684 (finetuned), it frequently degrades the primary classification performance of larger open-source models (32B–38B). For example, InternVL3-38B drops from 0.784 (zero-shot) to 0.767 (finetuned) on primary categories. This likely stems from overparameterization under data scarcity: our ~1,500 training samples create a capacity mismatch for larger models, leading to overfitting, whereas 7B–8B models achieve a more optimal bias-variance tradeoff.
Impact of Prompting Strategies
@@
Few-shot prompting yields modest boosts in primary-category performance for some models (e.g., GPT-4.1 improves from 0.796 to 0.813 accuracy), highlighting the value of in-context learning for broad tasks. However, in subcategory detection, few-shot gains are minimal or inconsistent. This suggests that for fine-grained tasks, a small set of exemplars may act as noise, anchoring the model to idiosyncratic patterns rather than clarifying the complex boundaries between subtypes. By contrast, supervised finetuning delivers substantial improvements across subcategory tasks, underscoring that direct model adaptation is the most effective strategy for specialized harmful content detection.
5.6 Per-Class Detection Performance
Commercial Models Open-Source Models
Claude-4 Gemini-2.5 GPT-4.1 InternVL3-38B Qwen2.5VL-32B
Primary Category ZS FS ZS FS ZS FS ZS FS FT ZS FS FT
Body 0.749 0.757 0.716 0.697 0.740 0.739 0.700 0.733 0.702 0.586 0.693 0.689
Food 0.866 0.856 0.829 0.809 0.807 0.840 0.832 0.842 0.845 0.780 0.854 0.829
Exercise 0.762 0.727 0.734 0.694 0.713 0.728 0.708 0.725 0.719 0.655 0.728 0.720
Supplements 0.922 0.928 0.892 0.882 0.897 0.924 0.898 0.928 0.890 0.870 0.882 0.901
Masculinity 0.867 0.869 0.828 0.771 0.832 0.847 0.793 0.818 0.549 0.735 0.844 0.654
Irrelevant 0.794 0.769 0.826 0.796 0.766 0.796 0.777 0.798 0.843 0.827 0.841 0.863
Table 3: Per-Class F1 Scores for Primary Category Classification by Model and Training Paradigm. ZS = Zero-Shot, FS = Few-Shot, FT = Finetuned. Highest scores per category in green, 2nd highest in blue.
Per-class performance analysis reveals systematic variation in detection difficulty that validates the importance of expert clinical knowledge (Table 3 for Task 1 and Table 12, Appendix E.2 for Task 2).
Content Explicitness Hierarchy
A clear hierarchy emerges: Supplement Abuse achieves the highest performance across all models (up to 0.928 F1) as harmful indicators like steroid mentions (“tren”, “stacking”) provide unambiguous lexical and visual signals. In contrast, Relationship to Body and Relationship to Exercise categories prove most challenging (F1 $\approx$ 0.70–0.76). These categories require distinguishing legitimate fitness content from harmful messaging along subjective boundaries, creating a dependency on expert knowledge that even large commercial models struggle to navigate without specific tuning.
Cultural Nuance and Data Density
Commercial models outperform open-source alternatives on culturally nuanced categories like Relationship to Masculinity, likely due to broader pre-training on sociological concepts. However, expert-labeled finetuning allows open-source models to yield results competitive with commercial systems in specific domains (e.g., InternVL3-38B matches Claude-Sonnet-4 on Supplements). Notably, finetuning effectiveness correlates directly with annotation density: while these models excel on data-rich Irrelevant content ($N=548$), they struggle significantly on sparse categories like Masculinity ($N=164$).
@@ 5.7 Severity Score Prediction
Claude‐Sonnet‐4 0.679 0.675
GPT‐4.1 0.690 0.690
Gemini‐2.5‐Flash 0.693 0.691
InternVL3‐8B 0.805 0.474
InternVL3‐38B 0.701 0.607
Qwen2.5‐VL‐7B 0.794 0.484
Qwen2.5‐VL‐32B 0.688 0.601
Table 4: Zero-shot Severity Prediction Metrics. MAE: Mean Absolute Error (lower is better); $\rho$: Spearman’s Rank Correlation (higher is better). Highest values in green, 2nd highest in blue.
To assess the ability of VLMs to predict harm severity, we conducted zero-shot prompting experiments across multiple models. Results shown in Table 4 reveal that models typically misestimate severity by approximately half a scale point, indicating reasonable but imperfect alignment with expert clinical judgment. Notably, severity prediction proves more challenging than categorical classification, with all models showing only moderate correlation with ground truth ratings. While these findings demonstrate baseline competence in severity estimation, substantial improvement is needed for reliable moderation deployment.
@@ 5.8 Error Analysis
Cross-model error analysis exposes fundamental challenges in characterizing pro-bigorexia content, particularly when visual signals conflict with textual context.
Qualitative Case Study
We illustrate these challenges through a representative error in an example video (Video_ID 030). The ground truth label is Relationship to Masculinity because the on-screen text explicitly ties specific body metrics (“170-190 lbs”, “1.6 shoulder-to-waist ratio”) to female desirability (“what women find”). While Claude-Sonnet-4 (zero-shot) correctly identifies this Masculinity context, both Gemini-2.5-Flash and GPT-4.1 (zero-shot) misclassify it as Relationship to Body. This error suggests that the misclassifying models overweight the visual modality, which features prominent shirtless posing, while underweighting the crucial textual overlay that reframes the physique display as a prerequisite for male social worth. This exemplifies the difficulty of distinguishing between pure body objectification and masculinity-driven dysmorphia without deep multimodal reasoning.
Systematic Misclassification
Broader quantitative analysis mirrors these qualitative findings. Models consistently misclassify exercise-related content as Irrelevant (0.061–0.143) or confuse Relationship to Body with Relationship to Exercise (0.102–0.173), reflecting annotator disagreements. Videos demonstrating workout sessions simultaneously touch on exercise, body display, dieting, lifestyle, and motivations that single-label classification cannot capture. Classification consistently confuses Hormone Therapy and Anabolic Steroids (0.118–0.176 bidirectional misclassification), a clinically critical distinction. Models also frequently misclassify Supplement Abuse as Relationship to Body (0.082–0.224), suggesting difficulty in recognizing harmful supplement messaging. This likely stems from creators using coded language (e.g., “tren”, “stack”), requiring domain expertise that general-purpose models lack. Additionally, models show systematic bias toward predicting Muscularity Self-objectification in subtype classification, with excessive exercise and toxic motivation content frequently misclassified into this category (0.147–0.471). These systematic errors highlight a fundamental challenge: effective pro-bigorexia detection demands not just multimodal capabilities, but clinical and social knowledge to navigate the blurred boundaries between fitness content and promoting unhealthy behaviors.
@@ 5.9.1 Input Features: Text vs Video
Gemini‑2.5‑Flash Audio 0.700 0.423 0.424
Audio All 0.669 0.471 0.478
@@
Table 5: Ablation results for task 1 using a single modality input features (text or video). Audio modality refers to the audio transcript of the video whose audio is classified as speech; Audio All modality refers to the audio transcript of all videos; Caption refers to the video description; OCR is text within images, and Text modality refers to Caption + OCR. Highest values in each column are in green, second‑best in cyan. Pm, Rm, and F1m refer to macro precision, macro recall, and macro F1 scores.
@@ 5.9.2 Number of Frames per Video
Table 6: Frame‑count ablation for GPT‑4.1 and InternVL3‑38B. Pm, Rm, and F1m are macro precision, macro recall, and macro F1. Highest values in each column are in green, second‑best in cyan.
@@ 6 Conclusion
We introduce BigTokDetect to bridge the gap between generic content moderation and the clinical nuances of muscle dysmorphia. While focused on this specific pathology, our work targets a broader class of multimodal mental health risks that evade traditional guardrails. Crucially, our results demonstrate that identifying these subtle behavioral harms relies less on model size and more on domain-specific adaptation; supplying high-quality, expert-annotated data for finetuning proves more effective than relying on massive generalist models. Future work may extend this via RLHF to integrate annotator disagreement as meaningful signal, preserving model utility while capturing the ambiguity inherent in clinical assessment that rigid training often discards. This functionality is particularly critical as short-form video platforms like TikTok increasingly shape youth culture and to combat the use of generative AI in health misinformation. Our framework provides reproducible methodology for the dynamic, behavior-aware moderation systems necessary to address the growing complexity of modern online harms.
@@
We did not track creator gender due to the privacy and ethical risks of visual inference. Since pro-bigorexia content typically skews male, while our “Irrelevant” control is likely gender-diverse, this demographic discrepancy presents a potential confounder for future work to address via privacy-preserving methods.
@@
This research addresses harmful content classification, raising important ethical considerations. To protect annotators reviewing potentially disturbing pro-bigorexia content, we recruited clinical experts, licensed psychologists, psychiatrists, and doctoral candidates specializing in eating disorders, plus computational social scientists with extensive experience in body image research. All annotators participated as voluntary research collaborators, fully informed of the project scope, with explicit rights to withdraw without penalty.
To strictly protect user privacy, we maintain strict protocols. While annotators accessed original content to ensure accurate clinical coding, we implemented a comprehensive manual de-identification protocol for the public data release. Four authors conducted a frame-by-frame review of the entire dataset. For every video, we annotated the specific on-screen coordinates of embedded usernames (@<username>), which can shift across frames, and redacted these regions using a white overlay box. Furthermore, we excised all platform-generated outro segments that display creator handles. This manual redaction, combined with the removal of user IDs and metadata, ensures that the stored videos and extracted frames contain no visible creator usernames or TikTok account identifiers. The full dataset will be made available only upon eligible research requests with ethics approval and signed data use agreements.
While supporting mental health research, we acknowledge risks of censorship against legitimate fitness communities. We focus on detection research rather than deployment recommendations. Our work is intended for research purposes and should not be used for clinical diagnosis without proper clinical oversight. The study protocol was reviewed and approved by the University of Southern California Institutional Review Board (IRB application number UP-23-00750).
We acknowledge the assistance of an AI language model in writing and editing portions of this manuscript. All AI-generated content was reviewed and edited by authors. Research design, data collection, analysis, and primary contributions remain entirely human work.
Acknowledgments
We are deeply grateful to our clinical collaborators at the University of California, Los Angeles Health, including Krista Tabuenca, Artha Gillis, Colleen McCord, Brandy Saccacio, Amanda Velkova, Chessa Kabiling, Jenny Nguyen, and Danielle Crowe, for generously volunteering their time and expertise to annotate this sensitive content, without whose clinical insights this work would not have been possible